Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, November 20, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 20, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/20/98 (Friday)

Segment 2B (Castaic to Sylmar): Report for 11 Nov 1998. At Romero Cyn, crews continued work installing electrical conduit and building frames for the vault site. Near Rye Cyn Rd, red concrete was poured over the completed electrical work supplying power to the valve site. At Feedmill Rd, directional bore crews have completed 1000' and are under the Santa Clara river. Crews on site at the river contained several small frac outs of drilling mud and continued monitoring the river for further leaks. Crews working south of Weldon Cyn completed tasks to the following mileposts working south to north. A splice box was set at EMP 72.81 although no fiber was placed in the ditch. Backfill completed 250' to EMP 72.85, asphalt completed 320' to EMP 72.85 while 200' of ditch was shaded to EMP 72.76. Coating crews have completed work to EMP 72.56, x-ray and welding to EMP 72.45, pipe crews have bent and laid pipe to EMP 72.44 while ditching is at EMP 72.23. Segment 2C (Templin Hwy to Castaic): General clean up of the right of way continued today and crews sent a cleaning pig through the pipe line to Weldon Cyn located within spread 2B.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 20 November 1998: 3A: Welding, x-ray, and coating crews worked on the bridge crossings at Pacoima and Tujunga Washes today, MP 78.63 and 81.69 respectively. Ditch crews also worked at Tujunga Wash. Backfill, padding, and fiber optic conduit crews focused on the north side of the Van Nuys bore area, MP 79.61. Slurry was also poured as part of that backfill operation. The bore at Tuxford Street continued, with crew pushing ~20 feet of casing pipe, bringing the total bore area to near 60 feet. That crew is being slowed by rocks encountered in the alignment. Fiber optic crews pulled cable between Tyburn Street and near the Glendale Freeway, at MP 94.67. 3B: No work occurred. 3C: No work occurred at the Hubbard bore. Final slurry was poured in San Fernando Road, MP 74.30. Work on the aqueduct cover occurred, MP 73.93. Crews continued to dig out the deep ditch in San Fernando Road, connecting Segments 2 and 3, to MP 73.87.

Segment 4B(El Segundo to Lynwood): The crew at 108 Th St. ESEMP 1.16 was installing the fiberoptic cable. The crew at 109 Th St. ESEMP 1.23 was doing electric work. The crew at east side of Crenshaw was doing electric work. The crew at Western ESEMP 5.56 was excavating for the fiberoptic conduit and they damaged the conduit while excavating. The crew at Sepulveda ROW ESEMP 10.78 and at Lynwood station EMP 0.0 were filling the pipe whit water for the Hydro test. The crew at El Segundo Station while backfilling damaged the off-set on Wednesday and they repaired it today.

Segment 5 (Lynwood to Long Beach) Report for 20 November 1998: The directional drill crew at Highway 91/Compton Creek was able to continue the ream despite the broken stem pipe that occurred yesterday. Progress is very slow. The frac-outs that are occurring are being contained and vacuumed so drilling mud is not entering Compton Creek. The bore crew at Wilmington refinery continues to push pipe through the casing.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page